OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of -3 and contains the point (4, -5). @love_jessika15

OpenStudy (anonymous):

\[y - \left( -5 \right) = 3 \left( x - 4 \right)\]

OpenStudy (anonymous):

\[y = 3x - 17\]

OpenStudy (anonymous):

\[-3x + y = -17\]

OpenStudy (anonymous):

or

OpenStudy (anonymous):

\[3x-y=17\]